OpenStudy (anonymous):

What's the volume of a sphere with a diameter of 34 inched?

OpenStudy (anonymous):

\[ V = \frac{4}{3} \pi r^3 = \frac{4}{3} \pi \left( \frac{d}{2} \right)^3 = \frac{4}{3} \pi \times 17^3 \]